Assume the following information:

You have $900,000 to invest

Current spot rate of Australian dollar (A$) is $0.62

180-day forward rate of the Australian dollar is $0.64

180-day interest rate in the U.S. is 3.5%

180-day interest rate in Australia is 3.0%

What is the return obtainable after 180 days from covered interest arbitrage?
Answer #1

An Arbitrage can be executed as described below:

- Borrow the Investment Corpus worth $ 900000 at the USD Interest Rate of 3.5 % for 180 days

- Borrowing creates a repayment liability worth (900000 x 1.0175) = $ 915750

- Convert the $ Investment Corpus into A$ at the current spot rate of $0.62/A $ to obtain 900000 / 0.62 = A$ 1451612.9

- Invest the converted A$ amount at the 180-Day Australian Interest Rate of 3.0 % to obtain an investment yield worth (1451612.9 x 1.015) = A $ 1473387.097

- Convert the investment yield into $ at the 180-day forward rate of $ 0.64 / A $ to obtain (1473387.097 x 0.64) = $ 942967.742

- Arbitrage Profit = 942967.742 - 915000 = $ 27217.742